# Short history of signed exceptional sequences

### v1

@@ 4.3. Higher dimensional cells of the picture space
To finish construction of the picture space, we need wide subcategories.
@@
The advantage of the picture space for $A_{n}$ is that it is very small and it is not to difficult to compute its homology. There is an obvious homomorphism from the Stasheff group $Sta_{n}=G(A_{n-1})$ to the unipotent group $U_{n}({\mathbb{Z}})$ given by sending generators $x_{ij}$ of $Sta_{n}$ to the elementary matrix $E_{ij}$ given by adding the $i$th column of the identity matrix $I_{n+1}$ to its $j$th column. In [19], we were motivated by the hope that the entire integral homology of the group $U_{n}({\mathbb{Z}})$ could be captured by copies of this construction.
@@ 5. Simplices to cubes
Given an hereditary algebra $\Lambda$ of finite representation type, its picture space is determined by the valued quiver $Q$ of $\Lambda$. The indecomposable $\Lambda$-modules are $M_{\alpha}$ where $\alpha\in\Phi^{+}(Q)$ runs over all positive roots of the root system of the Lie algebra associated to $Q$. For each $M_{\alpha}$ we have the semi-invariant domain $D(M_{\alpha})\subset{\mathbb{R}}^{n}$. We take intersection with the unit sphere $S^{n-1}$ and we get the semi-invariant picture
$\Sigma(Q)=\cup_{\alpha}D(M_{\alpha}).$
This set gives the n2n-2 skeleton of a triangulation of $S^{n-1}$. The n1n-1 simplices are labeled with the cluster-tilting object in the cluster category of $\Lambda$. These labels are distinct. For example, there are 14 triangles in $\Sigma(A_{3})\subset S^{2}$. We add a cone point (the origin in ${\mathbb{R}}^{n}$) and take $C\Sigma(Q)$. For $A_{3}$ this space is a union of 14 tetrahedra.
@@
    $D\Sigma(Q)$ has one oriented edge dual to each n2n-2 simplex $\lambda$ which is contained in the normally oriented $D(M_{\alpha})$ for a unique $\alpha$. This edge is labeled $e_{\alpha}$.
@@
    For each $k$-simplex $\rho$ in $\Sigma(Q)$, we have the dual cell $E(\rho)$ which is an nk1n-k-1 cell in $D\Sigma(Q)$. This dual cell $E(\rho)$ is labeled with the wide subcategory ${\mathcal{W}}(\rho)$ which is the class of all $M$ so that $\rho\subset D(M)$.
@@
    $D(\rho)$ is isomorphic to $X({\mathcal{W}}(\rho))$, the picture space of this wide subcategory.
The picture space is defined to be $C\Sigma(Q)$ with all dual cells on its surface with the same labels identified. Since the identifications are all on the surface, the interior of the $n$ remains a single $n$-cell which, when convenient, we view as a union of $n$-simplices, one for every cluster tilting object of $\Lambda$. We call the first the minimal cell structure of $X(Q)$, the second is the cubical cell structure of $X(Q)$.
@@
    The single vertex given by identifying all barycenters of all triangle in $\Sigma(A_{3})$. This is for the zero wide subcategory.
@@
    one 3-cell by definition of minimal cell structure. This is ${\mathcal{W}}=mod\text{-}\Lambda$
@@ 6. Cluster morphism category
Take cluster morphisms $S]:{\mathcal{W}}\to{\mathcal{W}}^{\prime}$ and $R]:{\mathcal{W}}^{\prime}\to{\mathcal{W}}^{\prime\prime}$. First, choose a simplex $\rho$ in $\Sigma(Q)$ so that ${\mathcal{W}}={\mathcal{W}}(\rho)$. To find one such $\sigma$, take the left Hom-Ext perpendicular category ${\mathcal{W}}_{0}=\,^{\perp_{01}}{\mathcal{W}}$. Choose a tilting module $T$ in ${\mathcal{W}}_{0}$, i.e., a cluster-tilting object with no shifted projectives. The $g$-vectors of the components of $T$ span a simplex $\sigma$ so that ${\mathcal{W}}(\sigma)=T^{\perp_{01}}={\mathcal{W}}$. Now compose the cluster morphisms
@@
This gives a cluster morphism $T^{\prime}]:mod\text{-}\Lambda\to{\mathcal{W}}^{\prime}$ where $T^{\prime}=T\oplus S^{\prime}$. The components of $T^{\prime}$ span another simplex $\tau$ containing $\sigma$ and $T^{\prime}]$ is the cluster morphism determined by $\tau$. This gives the following commuting diagram:
@@
This show that, given $\sigma$ so that ${\mathcal{W}}={\mathcal{W}}(\sigma)$ and a cluster morphism $S]:{\mathcal{W}}\to{\mathcal{W}}^{\prime}$, there is a unique simplex $\tau$ containing $\sigma$ so that the above diagram commutes where $[T]$ and $T^{\prime}]$ are the cluster morphisms determined by $\sigma$ and $\tau$.
@@ 7. Signed exceptional sequences
Lemma 7.1 (I-Todorov).
Every cluster morphism $[A]$ of rank $k$ has exactly $k!$ factorizations into a composition of $k$ morphisms of rank $1$, rank 1 morphisms being those given by a single object. These factorizations are in bijection with total orderings of the components of $A$.
@@
Theorem 7.4 (I-Todorov).
There is a bijection between (complete) signed exceptional sequences and ordered clusters.
We give some examples to illustrate this. We use semi-invariant pictures with chambers labeled with torsion classes. Recall that a torsion class $\mathcal{T}$} in $mod\text{-}\Lambda$ is a class of modules closed under quotients and extensions. Torsion classes form a poset under inclusion. This poset is a lattice in the finite case. The standard construction associates a torsion class ${\mathcal{T}}({\mathcal{U}})$ to each chamber $\mathcal{U}$}. We give an alternate formula.
@@
This formula defines a torsion class for any point $g\in{\mathbb{R}}^{n}$. Let ${\mathcal{S}}({\mathcal{U}})$ be the torsion class associated to the chamber $\mathcal{U}$} according to the formula above. For $\Lambda$ of finite representation type, this is independent of the choice of $g\in{\mathcal{U}}$ [27, Thm 1.8]. Also, ${\mathcal{S}}({\mathcal{U}})$ is different for different $\mathcal{U}$} by [27, Thm 2.6].
It is known that, when a semi-invariant wall $D(M)$ separates two chambers $\mathcal{U}$}, $\mathcal{U}$$^{\prime}$ with $\mathcal{U}$$^{\prime}$ on the positive side then the torsion class ${\mathcal{T}}({\mathcal{U}}^{\prime})$ contains the torsion class ${\mathcal{T}}({\mathcal{U}})$ and is the minimal such torsion class which also contains $M$ [2]. The class ${\mathcal{S}}({\mathcal{U}})$ has these properties. Indeed, it is shown in [27, Lemma 1.9] that $M$ lies in ${\mathcal{S}}({\mathcal{U}}^{\prime})$ but not in ${\mathcal{S}}({\mathcal{U}})$. In [27, Lemma1.11] it is shown that for any $X\in{\mathcal{S}}({\mathcal{U}}^{\prime})$ not in ${\mathcal{S}}({\mathcal{U}})$, there is an epimorphism $X\twoheadrightarrow M$ whose kernel is shown to lie in ${\mathcal{S}}({\mathcal{U}})$ in the proof of [27, Thm 2.1]. Therefore, ${\mathcal{S}}({\mathcal{U}}^{\prime})$ is the smallest torsion class containing ${\mathcal{S}}({\mathcal{U}})$ which also contains $M$. Since ${\mathcal{T}}({\mathcal{U}})$ and ${\mathcal{S}}({\mathcal{U}})$ have the same properties, they are equal. ∎
@@
In Figure 13 we show the semi-invariant picture with walls $D(M)$ labeled with $M$ in a box and torsion classes in ovals in some of the chambers. We will examine the signed exceptional sequences associated with the triangle labeled $T_{3}$. The cluster-tilting object at this simplex is $S_{1}\oplus S_{3}\oplus P_{1}$ since $a,b,c$ are the $g$-vectors of these modules. Figure 14 shows a blow up of this triangle with cube category showing the associated cluster morphisms. The cone of this triangle has cone point the origin in ${\mathbb{R}}^{3}$. The solid 3-ball is the union of these cubes which are the cones over the 14 triangles in $S^{2}$ shown in Figure 13.
@@
At the corners $a,b,c$ of the triangle there are wide subcategories ${\mathcal{W}},{\mathcal{W}}^{\prime},{\mathcal{W}}^{\prime\prime}$ respectively. ${\mathcal{W}}_{0}=\{S_{2}\}$ since it lies only on the wall with brick label $S_{2}$. The morphisms are defined up to sign:
${\mathcal{W}}^{\prime}=\{S_{1},S_{2},I_{2}\}=S_{3}^{\perp}$. So, $mod\text{-}\Lambda\to{\mathcal{W}}^{\prime}$ is $\pm S_{3}$.
${\mathcal{W}}_{0}=\{S_{2}\}={\mathcal{W}}^{\prime}\cap I_{2}^{\perp}$. So, ${\mathcal{W}}^{\prime}\to{\mathcal{W}}_{0}$ is $\pm I_{2}$.
The morphism ${\mathcal{W}}_{0}\to 0$ (at the center of the triangle) is $S_{2}[1]$ since the triangle labeled $T_{3}$ is on the negative side of the wall separating it from $T_{7}$ and $T_{3}\subset T_{7}$. The triangle is on the positive side of the other two walls. So, the other two morphism to the center of the triangle are positive $I_{2}$ and $S_{3}$. This makes ${\mathcal{W}}^{\prime}\to{\mathcal{W}}_{0}$ positive $I_{2}$.
${\mathcal{W}}^{\prime\prime}=\{S_{3},S_{2},P_{2}\}=P_{1}^{\perp}$. So, $mod\text{-}\Lambda\to{\mathcal{W}}^{\prime\prime}$ is $\pm P_{1}$.
${\mathcal{W}}_{0}=\{S_{2}\}={\mathcal{W}}^{\prime\prime}\cap S_{3}^{\perp}$. So ${\mathcal{W}}^{\prime\prime}\to{\mathcal{W}}_{0}$ is $\pm S_{3}$. But it is parallel to $S_{3}$ morphism to the center of the triangle. So, it must be positive.
@@
$(S_{2}[1],I_{2},S_{3})\simeq(S_{2}[1],S_{3},\pm P_{1})$
which differ by the braid move $\sigma_{2}^{-1}$ given by $\sigma_{2}^{-1}(S_{2}[1],I_{2},S_{3})=(S_{2}[1],S_{3},X)$ where, up to sign these are exceptional sequences (so $X=\pm P_{1}$) and $\operatorname{\underline{dim}}X$ is congruent to $\operatorname{\underline{dim}}I_{2}=(1,1,0)$ modulo $\operatorname{\underline{dim}}S_{3}=(0,0,1)$. So, $X=P_{1}$ with $\operatorname{\underline{dim}}X=(1,1,1)$. So, the sign of $P_{1}$ is positive.
The six signed exceptional sequences labeling the 6 paths from the origin of ${\mathbb{R}}^{3}$ to the center point of triangle labeled $T_{3}$ are all the factorizations of the cluster morphism $T_{3}:mod\text{-}\Lambda\to 0$. The other four signed exceptional sequences are:
$(I_{2},S_{1},P_{1})$
Since $S_{1}$ is not relatively projective, it cannot be negative. Also $\operatorname{\underline{dim}}S_{1}-\operatorname{\underline{dim}}S_{2}[1]=\operatorname{\underline{dim}}I_{2}$.
$(S_{3},-P_{2},P_{1})$
since $P_{2}$ fits in the exceptional sequence and $\operatorname{\underline{dim}}P_{2}[1]$ is congruent to $\operatorname{\underline{dim}}S_{2}[1]$ modulo $\operatorname{\underline{dim}}S_{3}$.
$(S_{3},P_{1},S_{1})$
$(I_{2},S_{3},S_{1})$
The last two are left to the reader to verify. Here the cluster corresponding to $T_{3}$ is $T_{3}=\{S_{1},S_{3},P_{1}\}$ since this is a cluster, but that was an accident since $T_{3}$ is a cluster. Usually you take the support tilting object contained in the torsion class.
@@
giving the signed exceptional sequence $(-S_{2},I_{2},S_{3})$. In red we have the following sequence of cluster morphisms where ${\mathcal{W}}^{\prime\prime}=\{S_{2},S_{3},P_{2}\}$:
@@
giving another signed exceptional sequence $(-S_{2},S_{3},P_{1})$.

### v2

@@ 4.3. Higher dimensional cells of the picture space
The construction of the picture space $X(Q)$ will be by induction (recursion) on the size of $Q$. The recursive construction requires wide subcategories.
@@
It is well-known [30] that, in finite type, wide subcategories $\mathcal{W}$} of $mod\text{-}KQ$ are isomorphic to $mod\text{-}KQ^{\prime}$ where $Q^{\prime}$ is another quiver of finite type which is smaller than $Q$. Thus, by induction, we have a picture space for $\mathcal{W}$} given by
$X({\mathcal{W}}):=X(Q^{\prime}).$
@@
The advantage of the picture space for $A_{n}$ is that it is very small and it is not too difficult to compute its homology. There is an obvious homomorphism from the Stasheff group $Sta_{n}=G(A_{n-1})$ to the unipotent group $U_{n}({\mathbb{Z}})$ given by sending generators $x_{ij}$ of $Sta_{n}$ to the elementary matrix $E_{ij}$ given by adding the $i$th column of the identity matrix $I_{n+1}$ to its $j$th column. In [21], we were motivated by the hope that the entire integral homology of the group $U_{n}({\mathbb{Z}})$ could be captured by copies of this construction.
@@ 5. Simplices to cubes
Given an hereditary algebra $\Lambda$ of finite representation type, its picture space is determined by the valued quiver $Q$ of $\Lambda$. The indecomposable $\Lambda$-modules are $M_{\alpha}$ where $\alpha\in\Phi^{+}(Q)$ runs over all positive roots of the root system of the Lie algebra associated to $Q$. For each $M_{\alpha}$ we have the semi-invariant domain $D(M_{\alpha})\subset{\mathbb{R}}^{n}$. We take the intersection with the unit sphere $S^{n-1}$ and we get the semi-invariant picture
$\Sigma(Q)=\cup_{\alpha}D(M_{\alpha})\cap S^{n-1}.$
This set gives the (n2)(n-2)-skeleton of a triangulation of $S^{n-1}$. The (n1)(n-1)-simplices of this triangulation are labeled with the cluster-tilting objects in the cluster category of $\Lambda$. These labels are distinct. For example, there are 14 triangles in $\Sigma(A_{3})\subset S^{2}$. Let $C\Sigma(Q)$ be the convex hull of $\Sigma(Q)$ in ${\mathbb{R}}^{n}$. This is the cone on $\Sigma(Q)$ with cone point the origin in ${\mathbb{R}}^{n}$. Each (n1)(n-1)-simplex in $\Sigma(Q)$ gives an $n$-simplex in $C\Sigma(Q)$. For $A_{3}$ this space $C\Sigma(A_{3})$ is a union of 14 tetrahedra.
@@
    $D\Sigma(Q)$ has one oriented edge dual to each n2n-2 simplex $\lambda$ which is contained in the normally oriented $D(M_{\alpha})$ for a unique $\alpha$. [The normal orientation of $D(M_{\alpha})$ is given by the vector $\alpha$ which is perpendicular to $D(M_{\alpha})$.] This edge is labeled $e_{\alpha}$.
@@
    For each $k$-simplex $\rho$ in $\Sigma(Q)$, we have the dual cell $E(\rho)$ which is an nk1n-k-1 cell in $D\Sigma(Q)$. This dual cell $E(\rho)$ is labeled with the wide subcategory ${\mathcal{W}}(\rho)$ which is the class of all indecomposable $\Lambda$-modules $M$ so that $\rho\subset D(M)$.
@@
    $E(\rho)$ is isomorphic to $X({\mathcal{W}}(\rho))$, the picture space of this wide subcategory.
The picture space is defined to be $C\Sigma(Q)$ with all dual cells on its surface with the same labels identified. Since the identifications are all on the boundary surface $S^{n-1}$ of the $n$-disk $D^{n}$, the interior of the $n$-disk remains a single $n$-cell. We call this the minimal cell structure of $X(Q)$. Sometimes it is convenient to view the $n$-disk as a union of $n$-simplices, one for every cluster tilting object of $\Lambda$. We call this the cubical cell structure of $X(Q)$.
@@
    The single vertex given by identifying all barycenters of all triangles in $\Sigma(A_{3})$. This is for the zero wide subcategory.
@@
    one 3-cell by definition of the minimal cell structure. This is ${\mathcal{W}}=mod\text{-}\Lambda$
@@ 6. Cluster morphism category
Take cluster morphisms $S]:{\mathcal{W}}\to{\mathcal{W}}^{\prime}$ and $R]:{\mathcal{W}}^{\prime}\to{\mathcal{W}}^{\prime\prime}$. First, choose a simplex $\rho$ in $\Sigma(Q)$ so that ${\mathcal{W}}={\mathcal{W}}(\rho)$. To find one such $\rho$, take the left Hom-Ext perpendicular category ${\mathcal{W}}_{0}=\,^{\perp_{01}}{\mathcal{W}}$. Choose a tilting module $T$ in ${\mathcal{W}}_{0}$, i.e., a cluster-tilting object with no shifted projectives. The $g$-vectors of the components of $T$ span a simplex $\rho$ so that ${\mathcal{W}}(\rho)=T^{\perp_{01}}={\mathcal{W}}$. Now compose the cluster morphisms
@@
This gives a cluster morphism $T^{\prime}]:mod\text{-}\Lambda\to{\mathcal{W}}^{\prime}$ where $T^{\prime}=T\oplus S^{\prime}$. The components of $T^{\prime}$ span another simplex $\tau$ containing $\rho$ and $T^{\prime}]$ is the cluster morphism determined by $\tau$. This gives the following commuting diagram:
@@
This show that, given $\rho$ so that ${\mathcal{W}}={\mathcal{W}}(\rho)$ and a cluster morphism $S]:{\mathcal{W}}\to{\mathcal{W}}^{\prime}$, there is a unique simplex $\tau$ containing $\rho$ so that the above diagram commutes where $[T]$ and $T^{\prime}]$ are the cluster morphisms determined by $\rho$ and $\tau$.
@@ 7. Signed exceptional sequences
Lemma 7.1.
[22] Every cluster morphism $[A]$ of rank $k$ has exactly $k!$ factorizations into a composition of $k$ morphisms of rank $1$, rank 1 morphisms being those given by a single object. These factorizations are in bijection with total orderings of the components of $A$.
@@
Theorem 7.4.
[22] There is a bijection between (complete) signed exceptional sequences and ordered clusters.
We give some examples to illustrate this. We use semi-invariant pictures with chambers labeled with torsion classes. By a chamber we mean a path component of the complement in ${\mathbb{R}}^{n}$ of the union of the set of walls. In this paper, there are only finitely many walls. So, the chambers are open sets and our definition agrees with standard definition [4]. Recall that a torsion class $\mathcal{T}$} in $mod\text{-}\Lambda$ is a class of modules closed under quotients and extensions. Torsion classes form a poset under inclusion. This poset is a lattice in the finite case. The standard construction associates a torsion class ${\mathcal{T}}({\mathcal{U}})$ to each chamber $\mathcal{U}$}. We give an alternate formula.
@@
The given formula defines a torsion class $\mathcal{T}$} for any point $g\in{\mathbb{R}}^{n}$.
  1. (1)
    Let $0\to A\to B\to C\to 0$ be a short exact sequence with $A,C\in{\mathcal{T}}$. Then $g\cdot\operatorname{\underline{dim}}B=g\cdot\operatorname{\underline{dim}}A+g\cdot\operatorname{\underline{dim}}C>0$. Also, any submodule $B^{\prime}$ of $B$ is an extension of a submodule $A^{\prime}$ of $A$ by a submodule $C^{\prime}$ of $C$. Indeed, $A^{\prime}=A\cap B^{\prime}$ and $C^{\prime}$ is the image of $B^{\prime}$ in $C$. By the Snake Lemma, we have an exact sequence $0\to A/A^{\prime}\to B/B^{\prime}\to C/C^{\prime}\to 0$. So,
    $g\cdot\operatorname{\underline{dim}}(B/B^{\prime})=g\cdot\operatorname{\underline{dim}}(A/A^{\prime})+g\cdot\operatorname{\underline{dim}}(C/C^{\prime})>0.$
    So $B/B^{\prime}\in{\mathcal{T}}$ and $\mathcal{T}$} is closed under extensions.
  2. (2)
    $\mathcal{T}$} is closed under taking quotients by definition since any quotient of a quotient of $X\in{\mathcal{T}}$ is a quotient of $X$.
Let ${\mathcal{S}}({\mathcal{U}})$ be the torsion class associated to the chamber $\mathcal{U}$} according to the formula above. For $\Lambda$ of finite representation type, this is independent of the choice of $g\in{\mathcal{U}}$ [29, Thm 1.8]. Also, ${\mathcal{S}}({\mathcal{U}})$ is different for different $\mathcal{U}$} by [29, Thm 2.6].
It is known that, when a semi-invariant wall $D(M)$ separates two chambers $\mathcal{U}$}, $\mathcal{U}$$^{\prime}$ with $\mathcal{U}$$^{\prime}$ on the positive side then the torsion class ${\mathcal{T}}({\mathcal{U}}^{\prime})$ contains the torsion class ${\mathcal{T}}({\mathcal{U}})$ and is the minimal such torsion class which also contains $M$ [2]. The class ${\mathcal{S}}({\mathcal{U}})$ has these properties. Indeed, it is shown in [29, Lemma 1.9] that $M$ lies in ${\mathcal{S}}({\mathcal{U}}^{\prime})$ but not in ${\mathcal{S}}({\mathcal{U}})$. In [29, Lemma 1.11] it is shown that for any $X\in{\mathcal{S}}({\mathcal{U}}^{\prime})$ not in ${\mathcal{S}}({\mathcal{U}})$, there is an epimorphism $X\twoheadrightarrow M$ whose kernel is shown to lie in ${\mathcal{S}}({\mathcal{U}})$ in the proof of [29, Thm 2.1]. Therefore, ${\mathcal{S}}({\mathcal{U}}^{\prime})$ is the smallest torsion class containing ${\mathcal{S}}({\mathcal{U}})$ which also contains $M$. Since ${\mathcal{T}}({\mathcal{U}})$ and ${\mathcal{S}}({\mathcal{U}})$ have the same properties, they are equal. ∎
@@
In Figure 13 we show the semi-invariant picture for the $A_{3}$ quiver $1\to 2\to 3$ with walls $D(M)$ labeled with $M$ in a box and torsion classes in ovals in some of the chambers. In the remainder of this paper we will examine the signed exceptional sequences associated with the triangle labelled $T_{3}$ in Figure 14 below both algebraically and geometrically. The cluster-tilting object at this simplex is $S_{1}\oplus S_{3}\oplus P_{1}$ since $a,b,c$ are the $g$-vectors of these modules. Figure 14 shows a blow up of this triangle with cube category showing the associated cluster morphisms. The cone of this triangle has cone point the origin in ${\mathbb{R}}^{3}$. The solid 3-ball is the union of these cubes which are the cones over the 14 triangles in $S^{2}$ shown in Figure 13.
@@
At the corners $a,b,c$ of the triangle there are wide subcategories ${\mathcal{W}},{\mathcal{W}}^{\prime},{\mathcal{W}}^{\prime\prime}$ respectively. These are wide subcategories of rank $2$. We also have three wide subcategories of rank $1$ which lie on the edges of the triangle. We look at ${\mathcal{W}}_{0}=\{S_{2}\}$. This is the rank 1 wide subcategory associated to the wall with brick label $S_{2}$. Since this wall has endpoints $b,c$ we have cluster morphisms ${\mathcal{W}}^{\prime}\to{\mathcal{W}}_{0}$ and ${\mathcal{W}}^{\prime\prime}\to{\mathcal{W}}_{0}$. We will determine how these morphisms are labeled. The morphisms are defined up to sign:
${\mathcal{W}}^{\prime}=\{S_{1},S_{2},I_{2}\}=S_{3}^{\perp}$. So, $mod\text{-}\Lambda\to{\mathcal{W}}^{\prime}$ is $S_{3}$ or $S_{3}[1]$.
${\mathcal{W}}_{0}=\{S_{2}\}={\mathcal{W}}^{\prime}\cap I_{2}^{\perp}$. So, ${\mathcal{W}}^{\prime}\to{\mathcal{W}}_{0}$ is $I_{2}$ or $I_{2}[1]$.
The morphism ${\mathcal{W}}_{0}\to 0$ (at the center of the triangle) is $S_{2}[1]$ since the triangle labeled $T_{3}$ is on the negative side of the wall separating it from $T_{7}$ and $T_{3}\subset T_{7}$. The triangle is on the positive side of the other two walls. So, the other two morphisms to the center of the triangle are positive $I_{2}$ and $S_{3}$. This makes ${\mathcal{W}}^{\prime}\to{\mathcal{W}}_{0}$ positive $I_{2}$.
${\mathcal{W}}^{\prime\prime}=\{S_{3},S_{2},P_{2}\}=P_{1}^{\perp}$. So, $mod\text{-}\Lambda\to{\mathcal{W}}^{\prime\prime}$ is $P_{1}$ or $P_{1}[1]$.
${\mathcal{W}}_{0}=\{S_{2}\}={\mathcal{W}}^{\prime\prime}\cap S_{3}^{\perp}$. So ${\mathcal{W}}^{\prime\prime}\to{\mathcal{W}}_{0}$ is $S_{3}$ or $S_{3}[1]$. But it is parallel to the morphism $S_{3}$ to the center of the triangle. So, it must be positive.
@@
$(S_{2}[1],I_{2},S_{3})\simeq(S_{2}[1],S_{3},P_{1}\,\text{or}\,P_{1}[1])$
which differ by the braid move $\sigma_{2}^{-1}$ given by $\sigma_{2}^{-1}(S_{2}[1],I_{2},S_{3})=(S_{2}[1],S_{3},X)$ where, up to sign these are exceptional sequences (so $X=P_{1}$ or $P_{1}[1]$) and $\operatorname{\underline{dim}}X$ is congruent to $\operatorname{\underline{dim}}I_{2}=(1,1,0)$ modulo $\operatorname{\underline{dim}}S_{3}=(0,0,1)$. So, $X=P_{1}$ with $\operatorname{\underline{dim}}X=(1,1,1)$. So, the sign of $P_{1}$ is positive.
The six signed exceptional sequences labeling the 6 paths from the origin of ${\mathbb{R}}^{3}$ to the center point of triangle labeled $T_{3}$ are all the factorizations of the cluster morphism $T_{3}:mod\text{-}\Lambda\to 0$. The first two, given above and in Figure 14 below are:
  1. (1)
    $(S_{2}[1],I_{2},S_{3})$
  2. (2)
    $(S_{2}[1],S_{3},P_{1})$
The other four signed exceptional sequences are:
  1. (3)
    $(I_{2},S_{1},S_{3})$, noting that, since $S_{1}$ is not relatively projective, it cannot be negative. Also $\operatorname{\underline{dim}}S_{1}=\operatorname{\underline{dim}}S_{2}[1]+\operatorname{\underline{dim}}I_{2}$. So $(I_{2},S_{1},S_{3})=\sigma_{1}^{-1}(S_{2}[1],I_{2},S_{3})$.
  2. (4)
    $(S_{3},P_{2}[1],P_{1})=\sigma_{1}^{-1}(S_{2}[1],S_{3},P_{1})$ since $P_{2}$ fits in the exceptional sequence and $\operatorname{\underline{dim}}P_{2}[1]$ is congruent to $\operatorname{\underline{dim}}S_{2}[1]$ modulo $\operatorname{\underline{dim}}S_{3}$.
  3. (5)
    $(S_{3},P_{1},S_{1})=\sigma_{2}^{-1}(S_{3},P_{2}[1],P_{1})$.
  4. (6)
    $(I_{2},S_{3},S_{1})=\sigma_{2}(I_{2},S_{1},S_{3})$.
The last two are left to the reader to verify. Here the cluster corresponding to $T_{3}$ is $T_{3}=\{S_{1},S_{3},P_{1}\}$ since this is a cluster, but that was an accident since the torsion class $T_{3}$ is a cluster. Usually you take the support tilting object contained in the torsion class.
@@
giving the signed exceptional sequence $(S_{2}[1],I_{2},S_{3})$. In red we have the following sequence of cluster morphisms where ${\mathcal{W}}^{\prime\prime}=\{S_{2},S_{3},P_{2}\}$:
@@
giving another signed exceptional sequence $(S_{2}[1],S_{3},P_{1})$.
8. Acknowledgments
The authors thank Kent Orr and Jerzy Weyman for their inspiration and collaboration in other works which lead to this work. We would not have discovered these concepts without them. The authors also thank the anonymous referee for their interest in this work and for their very thorough review which lead to many improvements in both exposition and content of the paper. Finally, the first author is grateful to the Simons Foundation for its generous support during the writing of this paper: Grant #686616.